User Tools

Site Tools


operations:handover

Handover notes

Please use this page to write up any notes on anything that needs to be passed on to the next observer or should go into the end of experiment message.

Notes should be placed below the horizontal line as a new blog entry.

Continuing problems should be copied to the Current Issues page.

**Template for Handover notes:**

Be sure to add any specific notes about the experiment if required, for example “Special back end configuration for x reason”. Put as much detail as you can - the more the better.

It is ESSENTIAL to put your name after EVERY entry. Copy and paste the template into the new handover notes

====== Experiment Name ======

**Hobart 12m**

Disk VSN: abc-123 | Data volume at beginning: 2.34 GB

   * 1300UT a comment from me (user)

**Katherine 12m**

Disk VSN: | Data volume at beginning:  GB

   * 

**Yarragadee 12m**

Disk VSN: | Data volume at beginning:  GB

   * Experiment started okay (JMc).

Format for entering a comment It is important to stick to this formula - there is a good reason we need this information, so please include it!

Time stamp in UT * Comment * (Your Name)

For Example:

20:00UT: Wind stows, Missed scans 012-3456 to 012-3457 (Jim)

20:50UT: Mark5 lost connection, had to restart Mark5 and reconfigure DBBC, called Jamie on call to fix, missed scans 012-3456 to 012-4567 (Jim)

Again - as much detail as you can. It is important to remember that these notes will be sent to the correlator, so please make sure they make sense to others.


AUM006

The AUM005 and AUM006 schedules have a problem with the wrap information - every scan has the “counter-clockwise” wrap selected.

aum006.skd has been edited to adjust this to all “neutral” wrap instead - Hopefully this will reduce the number of lost scans.

slogit has already been run for all stations - please do not re-download the SKD file.

Experiment Name

Hobart 12m

Disk VSN: Flexbuffhb | Data volume at beginning: N/A

Katherine 12m

Disk VSN: | Data volume at beginning: GB

  • SAR track during this experiment. Two scans removed: 231-2047 & 231-2051 (JMc)

Yarragadee 12m

Disk VSN: | Data volume at beginning: GB

  • SAR track during this experiment. Two scans removed: 231-2139 & 231-2141 and 231-2144 is shortened (JMc)
2018/08/19 04:47 · Warren Hankey · 0 Comments
AUM005

All antennas, scan 2 (003-066) waiting at limit. Source not risen yet? Scan missed. Turns out this happens a lot. Was this schedule scheduled for the correct day/time?

  grep -B 40 'ERROR fl' aum005ke.log | grep scan_name for ke,yg
  grep -B 40 'ERROR fl' aum005hb.log | grep "source="

to get list of missed scans.

Missed scans problem is a scheduling issue - all scans/all telescopes are in the counter-clockwise wrap (W in SKD file)

also, for HB12m, there is now a delaylive2 which will plot the maser to dbbc3 difference.

     on ops8: delaylive2 hb

Hobart 12m

Disk VSN: flexbuff | Data volume at beginning: - GB

  • 16:00 UT Experiment started (sort of) okay. See above.
  • Eremote control not working for hobart12. Using Dave's backup control and stream_log. Most of the time this worked better anyway.

Missed scans: 2018.230.16:00:50.00:source=0003-066,000340.29,-064017.4,1950.0,ccw 2018.230.16:02:54.00:source=2255-282,225522.47,-281425.6,1950.0,ccw 2018.230.16:14:04.00:source=0229+131,022902.51,130940.6,1950.0,ccw 2018.230.16:17:37.00:source=1759-396,175913.50,-394012.1,1950.0,ccw +++++++++++ many more

Katherine 12m

Disk VSN: | Data volume at beginning: 11622.083 GB

  • 16:00 UT Experiment started (sort of) okay. See above.
  • ES OVEN in maser vnc is now blue instead of green. Reads 0.9.

Missed scans due to source below limit: 2018.230.16:02:55.88:scan_name=230-1603,aum005,ke,92,92 2018.230.16:17:38.83:scan_name=230-1618,aum005,ke,200,200 2018.230.16:31:26.85:scan_name=230-1632,aum005,ke,71,71 2018.230.16:33:26.87:scan_name=230-1634,aum005,ke,48,48 +++++++ run bash from top for full list

Yarragadee 12m

Disk VSN: | Data volume at beginning: 1316.247 GB

  • 16:00 UT Experiment started (sort of) okay. See above.

Missed scans due to source below limit: 2018.230.16:14:00.03:scan_name=230-1615,aum005,yg,146,146 2018.230.16:17:37.08:scan_name=230-1618,aum005,yg,200,200 2018.230.16:31:29.03:scan_name=230-1632,aum005,yg,69,69 +++++ many more. Run bash from top.

2018/08/18 15:40 · Atifur R. Khan · 0 Comments
R4855

Katherine

Disk VSN: HAY-0010 | Data beginning: 0 Gb

  • 18:30 UT At beginning of experiment Jamie had to delete info off of the HAY module because it was full. Otherwise experiment started OK.
  • 06:13 (229) UT source 1243-160 above elevation limit? tracking errors for scan 229-0612b
  • 7:19 UT ALARM: Delay through DBBC has changed Monitor for stability and reset DBBC if it drifts. Reconfiguring the DBBC. Affected scans: 229-0731 - 229-0740 (Simin)
  • 8:07 UT Restarting DBBC. Affected scans: 229-0812 - 229-0815a (Simin)
  • 9:01 UT Drifting again. Called Warren and he suggest to shut down the DBBC to cool it down(Simin)
  • 9:20 UT Having problem with VNC so Warren called Mick to solve the problem (Simin)
  • 10:18 UT Reconfiguring the DBBC. Affected scans: 229-0900 - 229-1016 (Simin)
  • 16:30UT Maser VNC down (Arwin)
  • 1630UT Intermittent error “ WARNING: error ib -4 gpib device time-out on response c2” has been happening since start of experiment. Everything seems normal(Arwin)

Yarragadee

Disk VSN: JIVE-020 | Data beginning: 0Gb

  • 18:00 UT Field system unresponsive, and mk5 crashed, so experiment started 12 mins late. Afftected scans 228-1830 - 228-1838c.
  • 18:42 UT Experiment started OK.
  • 16:32UT: slow internet to Yg, sys monitor not working, neither is vnc to time yg. (Arwin)
2018/08/16 22:15 · Atifur R. Khan · 0 Comments
AUM004
Experiment Name

Hobart 12m

Disk: Flexbuffhb | Data volume at beginning: N/A

  • 22:55 - Experiment started after backend wiring corrected. (JMc)

Katherine 12m

Disk VSN: HOB+0124 | Data volume at beginning: 7723 GB

  • 22:55 - Experiment started late due to issues with Hobart12. Otherwise OK (JMc).
  • 07:03 - Formatter clock synced. scan missed 228-0705, 228-0707 (Atifur).
  • 07:40 - Field system froze. PCFS, MK5, DBBC all were reset. scan missed 228-0723 to 228-0908 (Atifur).
  • 10:00 - DBBC stopped responding. Rebooted. Schedule started @ 10:06 UT (Atifur).
  • 11:00 - Drift in delay. DBBC rebooted. (Atifur)
  • 11:40 - Drift in delay. DBBC rebooted. (Atifur)
  • 13:00 - Looked like it was going to start drifting again, but stabilised. (Katie)
  • After Experiment - data between ~0900-1510 recorded to another module, required for the subsequent R4 experiment. This had to be deleted in order to observe R4855 - no data will be available for correlation. Please make sure to check that the correct module is selected if the mark5/pcfs has been power cycled, before resuming the schedule. (JMc)

Yarragadee 12m

Disk VSN: HOB+0130 | Data volume at beginning: 7637 GB

  • 23:01 - Experiment started late due to issues with Hobart12. Otherwise OK (JMc).
2018/08/15 23:40 · Jamie McCallum · 0 Comments
CRDS97

The experiment started late. The CRDS97 vex file had not been made available on the ftp site and the operator had not recieved a distribution.

Hobart 26m

Disk VSN: HOB+0133/16000 | Data volume at beginning: 12750.32 GB

  • Experiment started extra late. Nominal start time was 226.17:57:21 however Hobart wind was very high (Lucas).
    • consistent wind stows - 17:57:21 to 19:08:49 UT
    • 19:11:26 to 19:42:44
    • 19:54:07 to current
  • Stow released at 07:59:08 UT (Prad).
  • 08:54:53 UT - error m5 -900 : already off and error m5 -906 mark5 return code 6: inconsistent or conflicting request. The recording halted as a consequence of the above error (Prad).
  • 09:28:47 UT - For the previous 28.7 minutes, there was a watchdog warning which prevented the log from getting updated.
  • 09:30:33 UT - The recording started.
  • Stow activated from 09:49:01 - 10:24:49 UT (Prad).
  • Recording started 10:25:46 UT (Prad).
  • Stow activated from 10:55:17 - 11:20:35 UT (Prad).

Katherine 12m

Disk VSN: IAAE-018/8000 | Data volume at beginning: 0.00 GB

  • Apart from late, experiment started okay (Lucas).
    • First scan no0325 at 17:57:21 (note that experiment first scan was no319; Lucas).
  • The scan no. 0478(0923-0928UT as seen in the summary file) didn't take place. The vex file showed that the Katherine is not scheduled for that scan (possibly owing to source being under horizon). An minor mistake in the summary file (Prad).

Yarragadee 12m

Disk VSN: HOB+0103/16000 | Data volume at beginning: 0.00 GB

  • Apart from late, experiment started okay (Lucas).
    • First scan no0325 at 17:57:21 (note that experiment first scan was no319; Lucas).
2018/08/14 22:58 · Lucas Hyland · 0 Comments
R1855

Katherine 12m

Disk VSN: OSOD+121 | Data volume at beginning: 5329.345 GB

  • 1700UT - Experiment started okay (Patrick).

Yarragadee 12m

Disk VSN: BKG+0136 | Data volume at beginning: 0.096 GB

  • 1700UT - Experiment started okay (Patrick).
2018/08/13 16:50 · Atifur R. Khan · 0 Comments
/home/www/auscope/opswiki/data/pages/operations/handover.txt · Last modified: 2018/07/26 22:26 by Lucas Hyland